### v3

# Reconstructing Hand-Held Objects in 3D from Images and Videos

Abstract
Objects manipulated by the hand (i.e., manipulanda) are particularly challenging to reconstruct from Internet videos. Not only does the hand occlude much of the object, but also the object is often only visible in a small number of image pixels. At the same time, two strong anchors emerge in this setting: (1) estimated 3D hands help disambiguate the location and scale of the object, and (2) the set of manipulanda is small relative to all possible objects. With these insights in mind, we present a scalable paradigm for hand-held object reconstruction that builds on recent breakthroughs in large language/vision models and 3D object datasets. Given a monocular RGB video, we aim to reconstruct hand-held object geometry in 3D, over time. In order to obtain the best performing single frame model, we first present MCC-Hand-Object (MCC-HO), which jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs. Subsequently, we prompt a text-to-3D generative model using GPT-4(V) to retrieve a 3D object model that matches the object in the image(s); we call this alignment Retrieval-Augmented Reconstruction (RAR). RAR provides unified object geometry across all frames, and the result is rigidly aligned with both the input images and 3D MCC-HO observations in a temporally consistent manner. Experiments demonstrate that our approach achieves state-of-the-art performance on lab and Internet image/video datasets. We make our code and models available on the project website: https://janehwu.github.io/mcc-ho
Figure 1: We present a scalable approach to hand-held object reconstruction from monocular RGB images or videos that is guided by object recognition and retrieval. Results demonstrate that our method is able to generate realistic object geometry that is also faithful to visual observations and consistent across frames. Please see the supplementary materials for evaluation of temporal consistency.
1 Introduction
This paper addresses the problem of recovering hand-object interactions in 3D from monocular RGB images and videos, focusing in particular on reconstructing hand-held object geometry. Given the tremendous advancements in object recognition and retrieval (e.g., generation), such as GPT-4(V) [51] and text-to-3D models [45, 57, 7, 43], a natural question that arises is whether reconstruction can benefit from leveraging these other branches of computer vision. Particularly for the challenging setting of reconstructing hand-held objects, such an approach may be essential to building up available training data to Internet scale. See Figure 1 and the supplementary videos.
Recovering hand-object interactions in 3D is an important problem in both computer vision and robotics. First, we aim to solve a particular instance of a classic computer vision problem (along with recognition and segmentation): 3D reconstruction. It is also a particularly difficult setting because the hand and object need to be reconstructed from a single RGB image, where the hand is articulated and the object can be small with respect to the image size and partially occluded by the hand. For videos, hand-object trajectories need to additionally be temporally consistent. Second, in robotics, particularly manipulation, there is a great need for big data. Unlike other areas such as natural language processing and computer vision, where big data is already available on the Internet, large-scale trajectories of robots manipulating objects are not readily available. There has been a longstanding aspiration to leverage trajectories of human hands interacting with objects [66]; in terms of videos, this data is readily available on the Internet. However, the significant challenge is how to extract 3D hand-object trajectories from those videos. To this end, our objective is to develop an automated technique for taking raw Internet video and extracting 3D hand-object trajectories in a model-free manner. Note that this version of the problem is distinct from model-based approaches that assume a known object shape [55, 100] or use multiview methods to reconstruct the object in 3D in a separate step [13, 31].
Our method consists of three stages: (1) learning-based hand-object reconstruction, (2) object model retrieval, and (3) rigid alignment over time of the retrieved model with network-inferred geometry. In the first stage, our transformer-based model, MCC-Hand-Object (MCC-HO), estimates hand-held object geometry from each RGB image and 3D hand. In the second stage, Retrieval-Augmented Reconstruction (RAR) is used to improve reconstruction quality via recognition and retrieval. Specifically, the hand-held object is recognized by GPT-4(V) and a 3D object is retrieved using a text-to-3D generative model (e.g., Genie [45]). In the third stage, the retrieved 3D object is rigidly aligned with the input image or video using a combination of DINOv2 [52] and inferred 3D cues from MCC-HO. Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [4, 2, 41]. To demonstrate the scalability of our approach, we additionally obtain pseudo ground truth 3D hands and objects for previously unlabeled videos in the 100DOH dataset [71].
Figure 2: Given an RGB video and estimated 3D hands, our method reconstructs hand-held object trajectories in three stages. In the first stage, MCC-HO is used to predict hand and object point clouds. In the second stage, a 3D model for the object is obtained using Retrieval-Augmented Reconstruction (RAR). In the third stage, the 3D object model is rigidly aligned with DINOv2 [52] image features and network-inferred geometry in a temporally consistent manner.
2 Related Work
Object reconstruction. Reconstructing objects from images is a longstanding problem in computer vision, and the bulk of research in this area has focused on scenarios where the object is fully or almost fully unoccluded in the input view(s). Structure from motion (SfM) [22, 70, 72, 76] and SLAM [3, 50, 73] remain robust solutions particularly for multiview reconstruction and camera parameter estimation. Learning-based object reconstruction methods typically represent 3D geometry as point clouds [60, 61], meshes [17, 86, 88], voxel grids [90, 83], or implicit surfaces [54, 47, 49, 75, 84, 89]. Following the release of Objaverse [10] and Objaverse-XL [11], a number of works have used this large-scale 3D object dataset to train monocular object reconstruction models [40, 39, 42, 27]. The CO3Dv2 dataset [67] contains objects from 50 MS-COCO categories [38] and has also been used to train monocular object reconstruction models [89]. In practice, we find that off-the-shelf object reconstruction methods do not perform well on images of hand-object interactions due to occlusion from the hand, larger object pose diversity, etc.
Hand reconstruction. Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [56, 37, 46, 48, 82, 98, 101, 53, 69, 59] or monocular video [30, 87, 92]. These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [68]. Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [81, 97]. We use HaMeR [56] in this work as it has achieved state-of-the-art performance on in-the-wild images.
Model-based hand-object pose estimation. When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [1, 2, 55, 14, 24, 25, 74, 78, 93, 91, 77, 62]. The estimated hands can then be used to rigidly align the object with images or videos [2, 55, 100] via inverse rendering [44, 64]. Due to the specific challenges arising from hand-object interactions, e.g. highly occluded geometry, optimization-based approaches are considered the current state-of-the-art in this area. RHO [2] studied the problem of recovering hand-object pose from in-the-wild images, whereas prior work was centered around datasets captured in lab settings. RHOV [55] extended RHO to video data by adding, among other terms, temporal and optical flow losses. Aiming to add more physics-inspired constraints, HandyPriors [100] leverages a differentiable contact module to minimize relative sliding between frames. G-HOP [96] guides optimization with a generative hand-object prior.
Model-free hand-object reconstruction. Reconstructing hand-held objects without corresponding 3D models poses additional challenges compared to general object reconstruction due to frequent occlusions by the hand and significantly less publicly available 3D data of hand-object interactions. Existing approaches are either tailored to reconstruction from a single image [94, 23, 5, 6, 99, 16, 8, 58], RGB video [28, 95, 21, 96], or RGB-D video [85]. Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved 3D object model.
Hand-object datasets. Datasets containing hand-object interactions are either collected in lab environments [1, 20, 4, 34, 41], through large-scale capture [18, 19], or from Internet videos [9, 71, 2]. A number of hand-held object reconstruction models [5, 6, 8] were trained on either the HO-3D [20] or DexYCB [4] datasets. Since the combined number of object instances in these two datasets is 31, it is difficult for such approaches to generalize to unseen objects. HOI4D [41] is currently the largest available dataset with 3D hand and object labels (also used in [95]). There is also a comparably larger amount of data containing videos of hand-object interactions without 3D labels. Epic Kitchens [9] contains 55 hours of egocentric video of people in their personal kitchens. Ego4D [18] and Ego-Exo4D [19] are also egocentric datasets, and 100 Days of Hands (100DOH) [71] contains 131 days of footage collected from Internet videos. Following the release of [71], MOW [2] and RHOV [55] labeled a subset of images and video clips from 100DOH, respectively, with 3D hands and objects.
3 Method Overview
Our approach to reconstructing hand-object interactions from monocular RGB video combines feedforward network inference (MCC-HO) with object model retrieval (RAR) and subsequent rigid alignment. See Figure 2. We build on the insight that the presence of a human hand provides grounding in terms of both the 3D location and scale of the object. Thus, as in prior work [94, 8], the estimated object pose and shape is conditioned on the estimated 3D hand. Applying a hand-centric approach has several advantages. Since monocular image/video reconstruction is particularly ill-posed with respect to depth and scale, 3D hands can be used to disambiguate the location and size of the object. Furthermore, the location of the 3D hand provides a coordinate system for estimating object geometry; in the case of rotating, symmetric objects, the hand can disambiguate the pose of the object.
Given a monocular RGB image or video, our method reconstructs hand-held objects in three stages. First, our transformer-based model, MCC-HO, estimates 3D hand-object geometry from single images (Section 4). Simultaneously, a 3D object model is automatically synthesized by a text-to-3D generative model [45] through a technique we coin Retrieval-Augmented Reconstruction (Section 5). Finally, the 3D object model is rigidly aligned in a temporally consistent manner (Section 6).
4 Estimating Hand-Object Geometry
In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [89] for the specific task of hand-object reconstruction. Our model, MCC-Hand-Object (MCC-HO), has to deal with two major challenges. First, MCC assumes RGB-D images as input, whereas our model needs to work with RGB input, since our ultimate goal is to reconstruct from Internet videos. Second, we have to find a way to cope with the significant domain shift between the training datasets. MCC was trained on the CO3D dataset [67], which consists of clean, 360-degree videos where each object is unoccluded and placed on a solid surface. For MCC-HO, our training data is a relatively small collection of in-the-wild and lab datasets of hand-object interactions, where the hand-held object is at least partially occluded in the vast majority of images.
The inputs to MCC-HO are a single RGB image and 3D hand geometry. Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM/SAM2) [32, 65], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image. Hand geometry is parameterized by the MANO hand model [68], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$. The MANO hand skeleton contains 21 joints. At training time, the ground truth 3D hand is used. At test time, HaMeR [56] is used to infer the hand geometry.
The inferred geometry is represented as a neural implicit function $\rho(x)$ modeling the hand and object jointly, such that for any 3D location $x$, $\rho$ returns the occupancy probability $\sigma(x)$, RGB color $c(x)$, and segmentation label $m(x)$:
$\rho(x)=(\sigma(x),c(x),m(x))$ (1)
where $\sigma(x)\in[0,1]$, $c(x)\in[0,1]^{3}$, and $m(x)\in\{0,1,2\}$ to indicating a background, hand, or object point. The scale of the object geometry is normalized with respect to the input 3D hand.
4.1 Architecture
MCC-HO has an encoder-decoder architecture that is composed of transformers [79] and outputs an implicit representation of hand-object geometry. The input images and 3D hand are passed to the encoder to compute a feature map that is a function of both image and hand embeddings. These features are then used to condition the decoder, which maps any 3D location to occupancy, color, and hand-object segmentation. During training, a set of query 3D points $Q$ are uniformly sampled within the bounds of a normalized volume. At test time, vertices of a voxel grid are passed as query points.
Encoder. The encoder takes as input a single RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a posed MANO hand mesh defined in camera world space with vertices $v_{h}\in\mathbb{R}^{778\times 3}$ and faces $f_{h}\in\mathbb{R}^{1538\times 3}$. The camera intrinsics are either known or consistent with the estimated 3D hand (e.g., the camera parameters used during HaMeR inference [56]). In order to jointly encode the image and hand geometry, the hand mesh is rasterized and per-pixel 3D points for the hand are sampled. For each pixel $(i,j)$ in the image, a ray is cast from the camera aperture to the pixel center to determine the first intersection point with the hand mesh at a point $P_{0}(i,j)$ in camera world space. If an intersection exists, the 3D location of each point is barycentrically interpolated from the intersecting camera world space triangle vertices, denoted $\{v_{h,1},v_{h,2},v_{h,3}\}$, such that:
$P_{0}(i,j)=\alpha_{1}v_{h,1}+\alpha_{2}v_{h,2}+\alpha_{3}v_{h,3}\in\mathbb{R}^{3},$ (2)
with barycentric weights $\alpha$. The points $P_{0}$ are used to define a normalized camera world space, such that the 3D coordinates passed to the network are $P=s(P_{0}-\mu(P_{0}))/\sigma(P_{0})$ and geometry is inferred in this same coordinate system. Importantly, $s$ is a constant scaling factor used to ensure the object geometry can be contained by the bounds of this hand-normalized sampling space.
The image $I$ and hand points $P$ are encoded into a single representation $R$, where:
$R\coloneqq f(E^{RGB}(I),E^{XYZ}(P))\in\mathbb{R}^{N^{enc}\times C},$ (3)
and $E^{RGB}$ and $E^{XYZ}$ are identical to the transformers proposed in MCC [89], though we define $P$ differently, i.e., without an input depth image. The image transformer $E^{RGB}$ has a Vision Transformer (ViT) architecture [12] with $16\times 16$ patch embeddings. The point transformer $E^{XYZ}$ uses a self-attention-based patch embedding design that differentiates between seen and unseen pixel points. In particular, for points in $P$ where no ray-mesh intersection was found, a special $C$-dimensional embedding is learned in lieu of embedding 3D points. The function $f$ concatenates the output of the two transformers, each of which has dimensions $N^{enc}\times C$. $N^{enc}$ is the number of tokens used in the transformers and $C$ is the output channel size.
Decoder. The decoder takes as input the encoder output, $R$, as well as a set of query points $Q\in\mathbb{R}^{N_{q}\times 3}$ (we choose $N_{q}=1024$). These inputs are passed to a transformer with architecture adopted from MCC and inspired by Masked Autoencoders (MAE) [26]. Each query token is passed to three output heads that infer binary occupancy $\sigma$, RGB color $c$, and hand-object segmentation $m$, such that:
$Decoder(R,Q)\coloneqq(\sigma(Q),c(Q),m(Q)).$ (4)
The decoder layers that output occupancy and color align with MCC, but MCC-HO has an additional linear layer that infers multiclass segmentation of each point into either background, hand, or object labels.
As discussed in further detail in MCC, a point-query-based decoder presents a number of advantages with respect to computational efficiency and scalability. Since the encoder is independent of the point queries, any number of points can be sampled using the same encoder structure and feature map size. The decoder architecture is relatively lightweight and also enables dynamic changes to the point sampling resolution and query size at training and/or test time.
Losses. The training loss is a combination of occupancy, color, and segmentation:
$\mathcal{L}=\mathcal{L}_{\sigma}+\mathcal{L}_{c}(\sigma_{gt})+\mathcal{L}_{m}$ (5)
The occupancy loss $\mathcal{L}_{\sigma}$ is a binary cross-entropy loss comparing the predicted and ground truth occupancies of each query point. The color loss $\mathcal{L}_{c}$ is a 256-way classification loss computed for query points that have ground truth occupancy of 1. The segmentation loss $\mathcal{L}_{m}$ is 3-way classification loss for each query point corresponding to background (empty space), hand, and object labels.
5 Retrieval-Augmented Reconstruction
Like any other deep learning model, how well MCC-HO is expected to perform is related to the amount of training data available, e.g., the size of available hand-object datasets. These are currently all quite small, orders of magnitude smaller than the “big data” on which computer vision foundation models have been trained. How might we find a way to exploit these other models to our end? Specifically, recognition of objects is very well developed [51], and large collections of 3D object models [10, 11] exist and have facilitated advancements in 3D generative modeling [45, 57, 7, 43]. Thus, in order to further improve upon network-inferred object geometry, we exploit these “big data” models by using GPT-4(V) [51] and a text-to-3D generative model [45] to automatically detect and “retrieve” a 3D object corresponding to the hand-held object in an image. We call this technique Retrieval-Augmented Reconstruction (RAR), somewhat in analogy to Retrieval-Augmented Generation in natural language processing [35].
Given a representative frame of the video, we prompt GPT-4(V) to provide a detailed description of the object the hand is holding. This text caption is passed to Genie, a text-to-3D model from Luma [45], to retrieve realistic 3D object geometry and appearance (Genie is also used in [36]). Since Genie outputs geometry normalized to fit inside a unit cube, the scale of the 3D model must be adjusted with respect to the input 3D hand.
5.1 Rigid Alignment
After a 3D object model is retrieved, it is rigidly aligned with the input visual data and MCC-HO predictions. We begin by considering the single image case. First, the object scale is estimated using the predicted MCC-HO object point cloud, which is a reasonable approximation because the prediction is grounded by the input 3D hand. If the MCC-HO point cloud is $X\in\mathbb{R}^{N\times 3}$ and a sampled point cloud of the retrieved object model is $Y\in\mathbb{R}^{N\times 3}$, then the scale is defined as $s=\sqrt{\lambda_{X}^{0}}/\sqrt{\lambda_{Y}^{0}}$ where $\lambda_{X}^{0}$ is the largest eigenvalue of $X^{T}X$ and $\lambda_{Y}^{0}$ is similarly defined for $Y$.
The remaining parameters to be optimized are object rotation $R$ and translation $T$. In order to (1) guarantee a globally optimal solution while (2) maintaining fixed compute time, we discretize the space of all possible parameter values to construct finite state spaces $\{\mathcal{R},\mathcal{T}\}$. The 3D rotation group, SO(3), is discretized using the unit cube subdivision strategy proposed in [33]. Translation is defined as an offset from the mean location of the MCC-HO point cloud $\mu(X)$, and the space is discretized by constructing a small voxel grid centered at $\mu(X)$. In order to avoid intractably large state space sizes (e.g., $\mathbb{R}^{3}\times\mathbb{R}^{3}$), we first optimize for rotation and then translation. For each parameter, we select the state that results in the lowest error using the sum of the following two metrics.
MCC-HO Chamfer Distance. The Chamfer Distance (CD) between the MCC-HO point cloud and a transformed object model is the usual 3D evaluation metric as defined in Equation 8. Note that when optimizing for rotation, the object model (centered at the origin) is first rotated and then translated to the mean location of the MCC-HO point cloud.
DINOv2 PCA Similarity. The MCC-HO point cloud gives us a reasonable estimate for object pose, but using this signal alone may not result in poses that align well with the input image (particularly for geometry with symmetries and in-the-wild objects). Thus, we additionally use DINOv2 [52] features to determine visual similarities between a transformed object model and the input image. For each rotation/translation state, the transformed object mesh is rendered using Pytorch3D [64]. A PCA basis is constructed using the DINOv2 features of the input image, and the first three PCA components are computed for both the rendered features $\mathcal{F}_{j}$ and input image features $\mathcal{F}_{0}$. We compute the cosine similarity between the two as
${\textstyle E_{DINO}=1-\frac{1}{2}\left(\frac{\text{PCA}(\mathcal{F}_{j})\cdot\text{PCA}(\mathcal{F}_{0})}{\max(||\text{PCA}(\mathcal{F}_{j})||_{2}\cdot||\text{PCA}(\mathcal{F}_{0})||_{2}),\epsilon)}+1\right)}$ (6)
where $\mathcal{F}_{j}$ and $\mathcal{F}_{0}$ are both masked by the input image object silhouette. When ground truth object masks are not available, we use SAM 2 [65]. See Figure 3 for rotation results using example images from 100DOH [71].
Figure 3: We compute a PCA basis of DINOv2 features using all frames masked by the object silhouettes (one frame and its first three PCA components are shown on the left side of each example). The first three components of this basis are used to determine the maximum likelihood Genie object rotation for each frame (shown on the right side).
6 From Images to Videos
In the previous section, we showed how given a single image, we select the rotation and translation of the 3D object model that minimizes error metrics based on CD and DINOv2 similarity. Now, if we have as input a video stream, we can further exploit the additional cue of temporal smoothness—we expect adjacent frames to have rotation and translation states that are close to each other.
Given a sequence of video frames with timesteps $t$, we jointly optimize for per-frame rotations $R(t)$ and subsequently per-frame translations $T(t)$. The typical approach for temporal alignment is to use gradient-based optimization  [55, 95], which is prone to issues with local minima and has unbounded computation time. Aiming for robustness and efficiency, we propose using the Viterbi algorithm [80, 15, 29] to solve for a global solution to rigid object alignment across all frames of a video sequence. The Viterbi algorithm is a dynamic programming approach to obtaining the maximum a posteriori probability of the most likely sequence of hidden states $Q=q_{1},q_{2},\ldots,q_{N}$ that results in a sequence of observed events $O=o_{1},o_{2},\ldots,o_{N}$. The hidden states $q_{t}$ can take on values from a finite set of parameters $\mathcal{Q}$ (where in our case, $\mathcal{Q}=\mathcal{R}$ or $\mathcal{Q}=\mathcal{T}$). Following the single image case, the observations $o_{t}$ are a combination of the MCC-HO inferred object point clouds and the DINOv2 image features.
Each cell of the Viterbi trellis, $v_{t}(j)$, is the probability of ending up in state $j$ after seeing the first $t$ observations and passing through the most probable state sequence $q_{1},\ldots,q_{t-1}$. Given the probability of being in every state at time $t-1$, the Viterbi probabilities are computed via:
$v_{t}(j)=\max_{i=1}^{|\mathcal{Q}|}\hskip 2.84526ptv_{t-1}(i)a_{ij}b_{j}(o_{t}),$ (7)
where $a_{ij}$ is the transition probability from the previous state $q_{i}$ to the current state $q_{j}$, and $b_{j}(o_{t})$ is the state observation likelihood given the current state $j$. In particular, $a_{ij}$ is Rodrigues’ rotation error when optimizing for rotations and Euclidean distance when optimizing for translations. The emission cost $b_{j}(o_{t})$ is the sum of the two metrics defined in the single image discussion, e.g., CD with respect to the MCC-HO point clouds and DINOv2 PCA similarity with respect to the input images.
7 Experiments
Section 7.1 details the datasets that were used for training and evaluation. Section 7.2 defines the evaluation metrics used for comparisons. We also quantitatively (Section 7.3) and qualitatively (Section 7.4) evaluate against existing work and find that our approach achieves state-of-the-art performance on all three datasets used [4, 2, 41]. Also in Section 7.4.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach applied to in-the-wild Internet videos from 100DOH [71] to demonstrate the feasibility and scalability of our automated model retrieval and fitting. Section 7.5 includes ablation studies for MCC-HO, RAR, and each component of our Viterbi temporal alignment algorithm.
DexYCB MOW HOI4D
F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
HO [23] 0.24 0.48 4.76 0.03 0.06 49.8 0.28 0.51 3.86
IHOI [94] - - - 0.13 0.24 23.1 0.42 0.70 2.7
HORSE [58] 0.20 0.41 7.16 0.08 0.20 27.2 0.21 0.41 7.68
MCC-HO (Ours) 0.36 0.60 3.74 0.15 0.31 15.2 0.52 0.78 1.36
Table 1: We compare our method, MCC-HO, to prior works on held-out test images from DexYCB, MOW, and HOI4D. Chamfer Distance (cm2) and F-score (5mm, 10mm) are reported.
Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
DiffHOI [89] 0.62 0.91 0.8
G-HOP [96] 0.76 0.97 0.4
MCC-HO + RAR (Ours) 0.91 0.98 0.28
Table 2: Quantitative evaluation of video-based hand-held object reconstruction methods on the HOI4D test dataset. Mean metrics are reported for consistency with prior work.
7.1 Datasets
We experiment on four different datasets for hand-object reconstruction.
HOI4D. The HOI4D dataset [41] consists of 4,000 lab videos of hands interacting with specified categories of common objects. We selected all sequences containing one of 6 hand-held object categories, and divide the sequences into training, validation and test sets. In total, we used 1142 training sequences, 6 validation sequences, and 12 test sequences consistent with DiffHOI [95]. For the validation set, we selected one video per category and favored object instances that appear the least frequently in the dataset.
DexYCB. The DexYCB dataset [4] contains 1,000 lab videos of hands interacting with 21 YCB objects. We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences.
MOW/RHOV. MOW [2] is a dataset of 512 images from 100 Days of Hands [71] labeled with 3D hands and objects. We use the same splits as [94] with 350 training images and 92 test images. Additionally, we use 110 videos from RHOV [55], each of which is a 2 second clip centered at the corresponding MOW frame. In the RHOV dataset, 19 videos overlap with the MOW test set and thus we do not use them during training.
100DOH. The 100 Days of Hands dataset [71] contains 131 days of footage collected from Internet videos showing hands interacting with objects. The vast majority of videos do not have associated 3D hand or object labels, so we evaluate how well our approach generalizes using examples from this unlabeled subset.
For DexYCB [4], MOW [2], and HOI4D [41], all results generated using our approach use the ground truth 3D hand provided by each dataset. For 100DOH [71] results (e.g., examples not contained in MOW), we use HaMeR [56] to estimate the 3D hand.
7.2 Evaluation Metrics
Since the focus of our approach is on the object manipulated by the hand, we quantitatively evaluate object reconstruction quality using standard 3D metrics: CD (cm2) and F-score reported at 5mm and 10mm. Both metrics are computed on point sets. The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [69] and HOI4D optimizes hand pose using 2D keypoint annotations. For each example, we sample 10k points on the surface of the ground truth mesh consistent with [94, 95] (no transformations are applied to the ground truth geometry). The MCC-HO predicted point clouds are supersampled/subsampled to obtain 10k points as well. In order to account for unknown ground truth camera parameters, ICP with scaling [64] is used to align all predicted geometry with the ground truth mesh. Median metrics are reported unless otherwise noted.
Chamfer Distance. The CD metric we use between two point sets $A,B\in\mathbb{R}^{N\times 3}$ is:
$CD(A,B)=\frac{1}{N}\sum_{a\in A}\min_{b\in B}||a-b||^{2}_{2}+\frac{1}{N}\sum_{b\in B}\min_{a\in A}||a-b||^{2}_{2},$ (8)
where the point sets should have the same size. This equation is consistent with prior work [5, 6, 94, 95]. We report CD in cm2 as in[5, 6].
F-score. The F-score at a given threshold $t$ is defined as a function of precision $P$ and recall $R$:
$F_{1}(X_{pd},X_{gt})=2\vspace{1mm}\frac{P(X_{pd},X_{gt},t)\cdot R(X_{pd},X_{gt},t)}{P(X_{pd},X_{gt},t)+R(X_{pd},X_{gt},t)}.$ (9)
7.3 Quantitative Evaluation
We quantitatively evaluate MCC-HO on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [23, 94, 58]. See Table 1. In order to sample points more densely in hand-object regions, the dimensions and granularity of the voxel grid used to query MCC-HO are determined from an initial coarse prediction using a default voxel grid. For DexYCB and HOI4D, we use every 5th frame of each test video sequence. The results for IHOI [94] and HO [23] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [95] papers; otherwise, we used available public repositories of prior work. Because AlignSDF [5] and gSDF [6] trained DexYCB-specific models, whereas we train a single model on multiple datasets, we do not compare to their published metrics. Our method achieves state-of-the-art performance for the Chamfer Distance and F-score metrics on all three datasets, implying that our technique is not overfitted to any particular dataset. We conjecture that increasing the number of unique object instances seen during training is important to model generalization. In addition, the pretrained MCC model provides a strong object prior for initialization.
We also compare to existing video-based approaches for RGB hand-held object reconstruction in Table 2, noting that the ground truth object is used for our RAR rigid alignment (a reasonable comparison as DiffHOI [95] and GHOP [96] train category-specific diffusion models using the ground truth HOI4D 3D objects).
7.4 Qualitative Evaluation
We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well retrieved object models using RAR can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks). Figure 4 shows predicted point clouds for a number of MOW test examples, as well as the result of rigidly aligning the ground truth CAD model with each point cloud. Since MCC-HO is trained with monocular images, the predicted RGB colors are only valid when viewed from the given camera. Because the object geometry is conditioned on the 3D hand, our approach tends to predict reasonable hand-object contact. Additionally, the object in contact with the hand is correctly identified by MCC-HO without input segmentation maps.
Figure 4: MCC-HO results for MOW test examples. The input image (top, left), network-inferred hand-object point cloud (top, right), rigid alignment of the ground truth mesh with the point cloud (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
Figure 5: Our approach applied to images in the 100 Days of Hands dataset [71]. For each image, we estimate 3D hands using HaMeR [56], inference object geometry using MCC-HO, retrieve a 3D model using RAR (e.g., GPT-4(V) and Genie), and rigidly align the 3D model with the inferred point cloud from MCC-HO.
7.4.1 Retrieval-Augmented Reconstruction
The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images and videos. Figure 5 shows our results for 9 images from 100DOH [71] that do not have 3D ground truth (e.g., not in MOW). Estimated 3D hands for each image are obtained using HaMeR [56]. At test time, our method does not require an input hand-object segmentation mask [2, 55], depth image [89], or hand joint locations [94, 8]. Please see the supplementary materials for additional results.
7.5 Ablation Studies
MCC Baseline. Because our model is built on top of MCC [89], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction. For MCC, we use the released model that was trained on the CO3Dv2 dataset [67] and apply an existing monocular depth estimator [63] to obtain depth maps (as MCC takes an RGB-D image as input). Table 3 compares 3D metrics between MCC and MCC-HO computed on the MOW test set. We find that using MCC-HO, which was trained on the specific task of hand-held object reconstruction, leads to performance improvements on test images.
Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC [89] 0.10 0.21 31.6
MCC-HO 0.15 0.31 15.2
Table 3: Analysis of how hand-held object reconstruction benefits from conditioning on the hand using the MOW test dataset.
RAR Success Rate. In order to quantify the success rate of RAR (e.g., whether a generated 3D object matches the input image), we use GPT-4(V) and Genie to automatically generate objects for all 512 images in the MOW dataset. Genie is able to generate an object 82% of the time, and otherwise returned an error. The results are included in the supplementary materials.
Video Alignment Components. We aim to understand the contributions of each cost term in our Viterbi algorithm, e.g., MCC-HO and DINOv2. See Table 4. In order to best assess the quality of rigid alignment, the ICP postprocess described in Section 7.2 is not applied. The results demonstrate that both terms contribute to the efficacy of our Viterbi algorithm; in practice, we observe that using both 2D and 3D cues is particularly important for objects that lie outside the training distribution of MCC-HO (DexYCB consists of commonplace YCB objects).
Viterbi Cost Terms F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC-HO 0.26 0.52 4.18
DINOv2 0.23 0.45 6.38
MCC-HO + DINOv2 0.27 0.52 4.10
Table 4: Analysis of how each of our Viterbi cost terms contributes to the final temporal alignment on the DexYCB test dataset.
8 Conclusion
In this paper, we present a novel paradigm for reconstructing hand-held objects that combines the respective benefits of model-free and model-based prediction. Our transformer-based model, MCC-HO, is trained to predict hand-object geometry given a single RGB image and estimated 3D hand. Experiments conducted using the DexYCB, MOW, and HOI4D datasets all demonstrate that MCC-HO achieves state-of-the-art performance in hand-held object reconstruction. We additionally present Retrieval-Augmented Reconstruction (RAR), an automatic method for object model retrieval that leverages recent advances in large language/vision models and 3D object datasets. These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH videos [71].
Limitations. As 3D generative models improve, RAR will naturally be able to retrieve more accurate and realistic geometry. Our approach also relies on the ability to obtain accurate hand and object masks which are tracked across frames. We expect that both MCC-HO and our rigid alignment algorithm will reap the benefits of any future image segmentation/tracking improvements.
Acknowledgements
We would like to thank Ilija Radosavovic for helpful discussions and feedback. This work was supported by ONR MURI N00014-21-1-2801. J. W. was supported by the NSF Mathematical Sciences Postdoctoral Fellowship and the UC President’s Postdoctoral Fellowship.